Status Report

NASA Space Station On-Orbit Status 3 Apr 2004

By SpaceRef Editor
April 3, 2004
Filed under , , ,
NASA Space Station On-Orbit Status 3 Apr 2004

All ISS systems continue to function nominally, except as noted previously or below.  Saturday — light-duty day for the crew, and their 166th day onboard ISS (168 days in space).

The crew performed the regular weekly 3-hr. station cleaning.  [“Uborka” includes removal of food waste products, cleaning of compartments with vacuum cleaner, wet cleaning of the Service Module (SM) dining table and other surfaces with “Fungistat” disinfectant and cleaning fan screens to avoid temperature rises.]

FE Alexander Kaleri deactivated the IK0501 gas analyzer (GA) in the SM and exchanged its BF carbon dioxide (CO2) filter assembly with a new unit from the FGB (replaced last: 2/20).  GA was reactivated and the spent BF stowed for return to Earth.  [IK0501 is an automated system for measuring CO2, O2, and H2O in the air as well as the flow rate of the gas being analyzed.]

Kaleri also performed the daily routine maintenance of the SOZh environmental control & life support systems in the SM and conducted the weekly inspection of the BRPK-1 water condensate separator.

Working from the Russian “job jar” task list, the FE downloaded measurements recorded during the recent (3/28-4/2) session of the Molniya-SM/LSO experiment from the LSO memory to HDD (hard-disk drive) on the EGE-2 laptop for return to Earth on Soyuz TMA-3 (7S).   [Objective of the unattended Molniya-SM session, similar to the French LSO experiment, was to record storm phenomena and other related events in the Earth’s equatorial regions.]

Sasha Kaleri had his weekly PFCs (private family conference), via VHF/home phone.

The crew also conducted the weekly planning conference with the ground, discussing next week’s “Look-Ahead Plan” (regularly prepared jointly by MCC-H and MCC-M planners), via S-band/audio.

Foale and Kaleri received kudos for yesterday’s work on the preliminary Progress 13P packing list, catching some errors that will help the ground to fine-tune the loading process.

Mike and Sasha worked out on their regular daily physical exercise program of 2.5 hrs on TVIS treadmill, CEVIS bike, RED exerciser and VELO cycle with load trainer.

TsUP/Moscow uplinked a preliminary list of equipment returning in the Descent Module (SA) of the Soyuz TMA 3/7S spacecraft on 4/29.  Crew was asked for a time estimate for stowing a number of bio kits and culture specimens that can be loaded only on undocking day.   [The 7S down-cargo identified currently totals 52.7 kg mass and 90.0 liters volume.  Return  cargo is classified as Station Status Monitoring, Russia, NASA, Japan, Holland, and ESA items.  NASA return items (1.88 kg) are a crew personal dosimeter, DST dual sorbent tubes (6) with reference kit, water samples, 24 FMK (formaldehyde monitoring kit) badges with reference kit, GSC (grab sample container) with April samples, and 16 local dosimeters.]

The Elektron O2 generator continues to operate nominally.  However, Russian controllers confirmed yesterday that its new Liquid Unit (BZh-7) is indeed leaking minutely, after its repressurization on 3/30.   [The hermetically sealed capsule inside the BZh is pressurized with nitrogen (N2) to 1.2 kg/cm^2, and the leak through the shell is reported to be so small that there is no cause for concern, unless leak rate increases dramatically.  There is enough N2 in the Elektron’s own N2 tank (8 liters @ 250 kg/cm^2 pressure) to feed the current leak for a longer period of time than the expected life cycle of the BZh.]

Data dumps of the two GNC MDMs (guidance, navigation & control computers) have verified that the RGA (rate gyro assembly) software patch loaded two days ago on both MDMs is working properly.   [The attitude control handover back from the Russian segment (RS) to the U.S. segment (USOS) was conducted with the new procedure that switches directly to momentum management without an attitude hold period for CMG (control moment gyro) desaturations.  This is accomplished by prepositioning the CMG gimbal angles to produce the desired momentum state to assume attitude control directly, without desats, in order to reduce stress on CMG-3.  The handover was delayed by one hour due to a mistake (wrong sign) made in the procedure to correct for a gimbal bias in the CMG-2 outer gimbal measurement (of +24.7 deg, known since 2001).  This bias had not been accounted for at first during the first handover on 3/27.]

The cause for the anomalous noise heard yesterday during the morning DPC (daily planning conference) has not been isolated yet.  [Past records indicate that at the time of noise ventilation fan #2 (VKYu-2) in the SM crew cabin cycled On/Off, and that the same fan also cycled On/Off at the time of the first occurrence of the “whumping” noise on 11/26/03.  Thruster activity (for attitude change) occurred yesterday over an hour later than the time of the noise.]

Evaluation of the video survey taken with the SSRMS (space station remote manipulator system) on 4/1 has shown that the H-Fixtures on the P1 truss are clear of MLI (multi-layer insulation), looking “great”, and that the CDRA (CO2 removal assembly) exhaust vent is clean (no ice buildup).

During the external video survey, Mike Foale traced the noise he reported during previous arm ops to the Lab camera’s Tilt motion (i.e., Pan is quiet, but he can clearly hear Tilt).

At yesterday’s SORR (Stage Operations Readiness Review) at JSC/Houston for 8Soyuz Launch/7Soyuz Return/Increment 9, the Go was given to the launch of Expedition 9 (Gennady Padalka, Mike Finke), with ESA taxi crewmember André Kuipers, on 8S/Soyuz TMA-4 on 4/18 (Eastern).  Padalka and Finke will return on 10/19 with 8S after a 181-day mission onboard ISS, while Kuipers returns with Expedition 8 (Mike Foale, Alexander Kaleri) in 7S/Soyuz TMA-3 on 4/29, after a docked mission duration of 9 days.  Other major events for Increment 9 are:

  • Undocking of Progress 13P — 5/18;
  • Launch of Progress 14P — 5/19;
  • Undocking of 14P — 7/21;
  • Launch of Progress 15P — 7/28;
  • and two Russian Orlan EVAs, on 7/22 (unique because no Soyuz docked on DC-1 & no Progress docked at SM aft end), and 8/24.

Systems that will be closely watched to make Increment 9 a success at this time include Elektron, TVIS treadmill, RED resistive exerciser, US EVA capability, minor Soyuz anomalies, some software changes planned for this year, and CMGs.

Weekly Science Update (Expedition Eight — 20th):

GASMAP:   Mike was thanked for his work with the GASMAP this Increment.

Human Research Facility/Workstation (HRF WS):   Continuing.

Advanced Ultrasound (ADUM):   No more activities scheduled for this Increment.

Hand Posture Analyzer (HPA:  Looking forward to operations in April.

In-Space Soldering Investigation (ISSE):   Nothing new.

Pore Formation and Mobility Investigation (PFMI):   Looking forward to activities on Monday, 4/5.
Space Acceleration Measurement System (SAMS):   Troubleshooting ICU (interface control unit) problem; was scheduled to work re-activation with the crew yesterday at 1:00pm EST.
Microgravity Acceleration Measurement System (MAMS):   Commanding on 4/2 was successful in recovering from a gimbal error.  Back to nominal operations.  Recovered from a RIC (rack interface controller) glitch this morning (4/3) with a HiRAP reset command.

Protein Crystal Growth-Single Locker Thermal Enclosure System (PCG-STES):  Behaving nominally.

Protein Crystal Growth Monitoring by Digital Holographic Microscope (PromISS):   Experiment has ended.

Investigating the Structure of Paramagnetic Aggregates from Colloidal Emulsions (InSPACE):    Planned.

Binary Colloidal Alloy Test-3 (BCAT-3):   The five files from the Friday (4/2) photo session are down and nearly ready for examination. The phase separation has been plotted over time and is
revealing interesting trends. The extra data will be extremely helpful.   Thanks to Mike for all the great images taken this past week.  Next BCAT-3 photo session is 4/5.

Renal Stone (RS):   The last in-flight collection is will be conducted 4/16-18.  Thanks to the crew for their continued support with daily pill ingestion.

Synchronized Position Hold, Engage, Reorient, Experimental Satellites (SHERES):   No more activities scheduled for this Increment.

Foot/Ground Reaction Forces during Space Flight (FOOT):   Looking forward to the last data collection session next week.

Materials ISS Experiment (MISSE):  In progress. Deployed outside. Nominal and collecting data.

Cellular Biotechnology Support Systems-Fluid Dynamics Investigation (CBOSS-FDI):   The ground received Mike’s video on FDI Tissue Culture runs, and the PI is looking at it.  Mike’s methods were called “great”.

Yeast Group Activation Packs (Yeast GAP):   Nothing new.

Earth Knowledge Acquired by Middle School Students (EarthKAM):  Nothing new. 

Earth Science Toward Exploration Research (ESTER):  No more activities scheduled for this increment.

Coarsening in Solid-Liquid Mixtures-2 (CSLM-2):  Planned.

Commercial Generic Bioprocessing Apparatus (CGBA):   CGBA continues to run nominally at 20C.

Miscible Fluids in Microgravity (MFMG):    Four runs are remaining; they will be performed on Increment 9.

Educational Payload Operations (EPO):   No more activities scheduled for this increment.

Crew Earth Observations (CEO):  Highest kudos for the crew for their excellent response to last week’s dynamic event request for the first-ever documented hurricane in the South Atlantic Basin.  Their fantastic 21-frame sequence from this past Saturday morning (3/27) has been requested by investigators at the National Hurricane Center and was posted on Earth Observatory.  Downlinked imagery continues to arrive at a prodigious rate, much faster than the ground can thoroughly review at this time.  It was noted that most of it is either panoramic or oblique and investigators assume that this is due both attitude and window constraints.  Camera times appear to be fine.  Thanks to the crew for their continued good efforts for CEO.

For CEO, spring weather conditions in the Northern Hemisphere will offer numerous, important visual changes on the Earth including snow and ice melt, vegetation greenup, agricultural activity, flooding and severe weather phenomenon.  Today’s optional CEO targets, limited in XPOP attitude by flight rule constraints on the use of the science window, which is available for only ~1/4 of each orbit when not facing forward (in “ram”), were Delhi, India (ISS pass had fair weather and excellent illumination.  Looking just right of track for the Indian capital city), Sobat Fans, Sudan (the White Nile passes through a large basin in this region that contains the Sudd Swamp and the Machar Marshes.  Outflow from the Ethiopian Plateau to the East has produced a complex of alluvial fans.  Good nadir mapping passes are needed for further study of this remote site), Logone Basin Fans, Chad (the complex of inland basins combined with long-term climate changes have produced numerous, poorly understood megafan series in this region.  As ISS tracked across this target area, the crew was to try for a nadir mapping sequence to begin baselining the character and distribution of these features), Cairo, Egypt (ISS had a nadir pass over the Egyptian capital), and Bamako, Mali (weather is expected to improve for this nadir pass over the Mali capital, situated on the upper Niger River).

CEO images can be viewed at the websites.

See also the website “Space Station Challenge” at

ISS Orbit  (as of this morning, 11:24am EST [= epoch]):

  • Mean altitude — 365.2 km
  • Apogee — 371.856
  • Perigee — 358.591m
  • Period — 91.85 min.
  • Inclination (to Equator) —  51.6294 deg
  • Eccentricity — 0.0009835
  • Orbits per 24-hr. day — 15.68
  • Mean altitude loss last 24 hours — 180 m
  • Revolutions since FGB/Zarya launch (Nov. ’98) — 30662

For more on ISS orbit and worldwide ISS naked-eye visibility dates/times, see

SpaceRef staff editor.